OpenStudy (anonymous):

ok i totally had a brain fart im in geometry and i forgot how to subtract fractions can someone help please?? 13 3/4 - 6 5/8

ganeshie8 (ganeshie8):

mixed fractions

ganeshie8 (ganeshie8):

subtract 13-6, then subtract 3/4 - 5/8

ganeshie8 (ganeshie8):

\(\huge 13 \frac{3}{4} - 6\frac{5}{8}\)

ganeshie8 (ganeshie8):

thats the given expression right?

OpenStudy (anonymous):

oh sorry but yes it is

OpenStudy (anonymous):

so with the 3/4 and the 5/8 would i find the common denominator than subtract??

ganeshie8 (ganeshie8):

thats right!!

OpenStudy (anonymous):

so 7 and 1/8???

ganeshie8 (ganeshie8):

\(\huge 13 - 6 + \frac{3}{4} - \frac{5}{8}\)

ganeshie8 (ganeshie8):

correct!

ganeshie8 (ganeshie8):

\(\huge 7+\frac{1}{8} = 7\frac{1}{8}\)

OpenStudy (anonymous):

ok thank you sooo much for helping me i cant believe i forgot that!!!!